ABC sells equipment to XYZ for $8,265. XYZ will pay $10,000 in two years. What is the implicit interest rate in this transaction?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

10,000=8,265*(1+r/100)^2

(10,000/8,265)^(1/2)=(1+r/100)

(1+r/100)=1.1

r=1.1-1

=10%(Approx).